# DeepAAT: Deep Automated Aerial Triangulation for Fast UAV-based Mapping

### 02 Feb 2024 v1

@@ V-D Comparison
ESFM [28]: A neural network architecture is proposed, which takes track points as input in the form of a matrix. It can simultaneously predict camera pose and scene points, and use reprojection error as the loss function.
Colmap [8]: A state-of-the-art open-source incremental SfM pipeline library, widely used in pose estimation and scene reconstruction tasks. Colmap provides both UI interface and command line running mode, making it easy to operate and has good reconstruction results.
OpenMVG [46]: Provides both incremental SfM and global SfM implementations, with global SfM being the current SOTA in the open-source library. By using the command line, step-by-step SfM can be easily implemented.
Because the output of ESFM and the proposed method is the result before BA, we directly compared the predicted results of the two networks, including reprojection error and rotation error. The results are shown in Tab.IV, and some prediction scenarios are shown in Fig.8. To make the learning space of ESFM more consistent, we standardize all input scenes, so that the network learns the relative positions of all cameras relative to the first camera in the scene. When comparing the proposed DeepAAT with Colmap and OpenMVG, we directly compared the final reconstruction results after BA, including reprojection error, final scene points, and time consumption. The results are shown in Tab.V, and the predicted scenes are shown in Fig.9.
TABLE IV: Comparison of ESFM and the proposed DeepAAT
@@
From Tab.IV, it can be seen that the prediction results of the proposed DeepAAT are much smaller than ESFM in terms of reprojection error, position error, and rotation error. From the four comparative scenarios in Fig.8, it can be seen that the scenarios predicted by the proposed DeepAAT are all correct, but the prediction results of ESFM are relatively chaotic, resulting in the inability to obtain correct results through subsequent global BA. The main reasons are as follows:
(1) DeepAAT integrates GPS information as a prior, significantly enhancing the network’s spatial awareness and location perception. Crucially, DeepAAT predicts the offset in camera position, which is a relative measure to GPS coordinates, rather than attempting to directly ascertain the precise location of each camera.
(2) Unlike ESFM, DeepAAT incorporates a global consistency-based outlier rejection module, which effectively eliminates erroneous matching points that persist even after geometric verification. As a result, the prediction outcomes produced by DeepAAT are considerably more refined and cleaner. In contrast, ESFM lacks a denoising feature, and the presence of noise points in its framework can adversely affect the network’s ability to accurately learn and represent the correct scene.
TABLE V: Comparison results between DeepAAT and traditional algorithms
@@
Table V reveals that DeepAAT outperforms all other methods in terms of average reprojection error and average time consumption, with both indicators surpassing those of Colmap and OpenMVG. Its most striking advantage lies in time efficiency, where DeepAAT significantly outpaces the comparison methods across all test scenarios. Specifically, DeepAAT’s average reconstruction efficiency is 453 times greater than Colmap, 580 times that of OpenMVG Incremental, and 28 times that of OpenMVG Global. This suggests that the proposed network substantially enhances the efficiency of AAT reconstruction while maintaining scene integrity. Concurrently, it also improves reconstruction accuracy to a certain degree. As indicated in Table V, OpenMVG Incremental failed to reconstruct scene 8_127, likely due to the stringent requirements of the incremental SfM algorithm on initial image pair selection and the relative instability of the OpenMVG Incremental algorithm.
V-E Ablation Study
To test the impact of the core modules proposed in DeepAAT, the following ablation experiments were conducted: 1⃝ The encoding order of GPS $L_{G}$ and the projection matrix $L_{prj}$ is set to 2; 2⃝ The encoding order of GPS $L_{G}$ and the projection matrix $L_{prj}$ is set to 4, which is the setting used in this article; 3⃝ The encoding order of GPS $L_{G}$ and the projection matrix $L_{prj}$ is set to 6; 4⃝ Remove the spatial spectral feature aggregation module, similar to ESFM, using only the projection matrix $\mathbf{W}_{prj}$ as the network input; 5⃝ Remove the global consistency-based outlier rejecting module. Here, the predicted pose of the network is directly used to triangulate all matching track points during global triangulation. These experimental data are the average of 10 test data results, which are shown in Tab.VI. From Tab.VI, it can be seen that:
TABLE VI: Results of ablation study
@@ V-E Ablation Study
(1) Encoding order $L$ of GPS information and the projection matrix in the spatial-spectral feature aggregation module has little impact on the experimental results, indicating that as long as the encoding order is set within a reasonable range, good experimental results can be achieved.
@@
(3) Upon removal of the global consistency-based outlier rejecting module, there is an increase in the reprojection error. As shown in Fig.9, the noise feature points generated by triangulation increases significantly. The outliers included in the scene can have a negative impact on subsequent BA and easily lead to local optimum. For example, for scene 8_127, the final result after BA is shown in Fig.10. It can be seen that there is a camera (marked with a green dashed box) whose pose has been optimized erroneously due to the influence of outliers.
@@
V-F Generalization experiments on different sizes of input images
The proposed network does not require a fixed number of images as input. Therefore, to test the generalization of the network in scenarios with inconsistent camera numbers compared to the training sample, we apply the model trained on scene from 100-130 images to predict scenes with 30-50 cameras and 400-430 cameras. The predicted results are shown in Tab.VII and Tab.VIII.
TABLE VII: Results of outlier rejecting for 100-130 camera models in different numbers of image prediction tasks
@@ V-F Generalization experiments on different sizes of input images
TABLE VIII: Pose prediction results of the 100-130 camera trained model in different numbers of camera prediction tasks
@@
The prediction results demonstrate that the proposed DeepAAT is versatile, excelling not only in scenes with a similar number of images but also in scenarios with significantly more or fewer images. As indicated in Tab.VII, the network trained on scenes with 100-130 images shows a decline in accuracy, precision, recall, and F1 score when applied to 30-50 image scenes. Conversely, its performance metrics improve in 400-430 image scenes. This improvement could be attributed to the longer average track length of matching points in scenes with more cameras, increasing the likelihood of points being classified as inliers and reducing false negatives. This explains the high recall of 0.993 in scenes with 400-430 cameras. Regarding prediction time, the network demonstrates a notable efficiency: despite a rapid increase in the number of cameras, the time required for network prediction increases only marginally. This efficiency is a significant advantage in practical applications. In traditional AAT algorithms, both incremental and global, time consumption escalates quickly with an increasing number of scene images, a trend particularly pronounced in incremental SfM. Thus, our network’s time-saving benefits become more pronounced with larger sets of scene images.
From Tab.VIII, we observe that the network, when trained on scenes with 100-130 images, exhibits a slight increase in average reprojection error, position error, and rotation error while predicting scenes with 30-50 images and those with 400-430 images. However, in scenarios such as 1_420 and 5_404, the network’s predicted position errors surpass the initial position errors. Nevertheless, following global BA, all scenes achieve accurate reconstruction results, as illustrated in Fig.11. This outcome not only underscores the network’s effective scene initialization capabilities but also highlights its robust reconstruction prowess.
@@ VI Conclusion
Biography Section
[Uncaptioned image] Zequan Che received the B.S. degree and M.S. degree in GIS from Wuhan University, Wuhan, China, in 2020 and 2023, and now he is a researcher in Guangzhou Urban Planning & Design Survey Research Institute, Guangzhou, China. His research interests include 3D reconstruction, UAV mapping, point cloud data processing, and so on.
[Uncaptioned image] Jianping Li received the B.S. degree in GIS and the Ph.D. degree in photogrammetry and remote sensing from Wuhan University, Wuhan, China, in 2015 and 2021, respectively. His research interests include 3D sensing system integration, UAV/UGV mapping, robot perception, and point cloud data processing.
[Uncaptioned image] Qusheng Li is currently with Guangxi Zhuang Autonomous Region Remote Sensing Institute of Nature Resources. He mainly engaged in surveying and mapping, geographical information, and 3D reconstruction. He is a National Surveying and Mapping Technology Expert.
[Uncaptioned image] Bisheng Yang received the B.S. degree in engineering survey, the M.S. degree, and the Ph.D. degree in photogrammetry and remote sensing from Wuhan University, China, in 1996, 1999, and 2002, respectively. From 2002 to 2006, he held a post-doctoral position at the University of Zurich, Switzerland. Since 2007, he has been a Professor with the State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing (LIESMARS), Wuhan University, where he is currently the Director of LIESMARS. His main research interests comprise 3-D geographic information systems, urban modeling, and digital city. He was a Guest Editor of the ISPRS Journal of Photogrammetry and Remote Sensing, and Computers & Geosciences.
[Uncaptioned image] Zhen Dong received the B.E. and Ph.D. degrees in remote sensing and photogrammetry from Wuhan University, in 2011 and 2018, respectively. He is currently a Professor of the State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing (LIESMARS), Wuhan University. His research interest lies in the field of 3D Computer Vision, particularly including 3D reconstruction, scene understanding, point cloud processing as well as their applications in intelligent transportation system, digital twin cities, urban sustainable development and robotics.

### 07 Apr 2024 v2

@@ 5.4 Comparison
ESFM (Moran et al., 2021): A neural network architecture is proposed, which takes track points as input in the form of a matrix. It can simultaneously predict camera pose and scene points, and use reprojection error as the loss function.
Colmap (Schonberger and Frahm, 2016): A state-of-the-art open-source incremental SfM pipeline library, widely used in pose estimation and scene reconstruction tasks. Colmap provides both UI interface and command line running mode, making it easy to operate and has good reconstruction results.
OpenMVG (Moulon et al., 2013a): Provides both incremental SfM and global SfM implementations, with global SfM being the current SOTA in the open-source library. By using the command line, step-by-step SfM can be easily implemented.
Because the output of ESFM and the proposed method is the result before BA, we directly compared the predicted results of the two networks, including reprojection error and rotation error. The results are shown in Tab.4, and some prediction scenarios are shown in Fig.8. To ensure consistency in the learning space of ESFM, we standardized all input scenes, allowing the network to learn the relative positions of all cameras with respect to the first camera in the scene. When comparing the proposed DeepAAT with Colmap and OpenMVG, we directly compared the final reconstruction results after BA, including reprojection error, final scene points, and time consumption. The results are shown in Tab.5, and the predicted scenes are shown in Fig.9.
Table 4: Comparison of ESFM and the proposed DeepAAT
@@
Tab.4, indicates that the prediction results of the proposed DeepAAT are much smaller than ESFM in terms of reprojection error, position error, and rotation error. From the four comparative scenarios in Fig.8, it can be seen that the scenarios predicted by the proposed DeepAAT are all correct, but the prediction results of ESFM are relatively chaotic, resulting in the inability to obtain correct results through subsequent global BA. The main reasons are as follows:
(1) By integrating GPS information as prior knowledge, DeepAAT significantly enhances its spatial awareness and perception of location. Crucially, DeepAAT predicts the offset in camera position, which is a relative measure to GPS coordinates, rather than attempting to directly ascertain the precise location of each camera.
(2) In contrast to ESFM, DeepAAT incorporates a global consistency-based outlier rejection module, which effectively eliminates erroneous matching points that persist even after geometric verification. As a result, the prediction outcomes produced by DeepAAT are considerably more refined and cleaner. In contrast, ESFM lacks a denoising feature, and the presence of noise points in its framework can adversely affect the network’s ability to accurately learn and represent the correct scene.
Table 5: Comparison results between DeepAAT and traditional algorithms
@@
Table 5 reveals that DeepAAT outperforms all other methods in terms of average reprojection error and average time consumption, with both indicators surpassing those of Colmap and OpenMVG. Its most striking advantage lies in time efficiency, as DeepAAT significantly outpaces the comparison methods across all test scenarios. Specifically, DeepAAT’s average reconstruction efficiency is 453 times greater than Colmap, 580 times that of OpenMVG Incremental, and 28 times that of OpenMVG Global. This suggests that the proposed network substantially enhances the efficiency of AAT reconstruction while maintaining scene integrity. Concurrently, it also enhances reconstruction accuracy to a certain extent. As indicated in Table 5, OpenMVG Incremental failed to reconstruct scene 8_127, likely due to the stringent requirements of the incremental SfM algorithm on initial image pair selection and the relative instability of the OpenMVG Incremental algorithm.
5.5 Ablation Study
To test the impact of the core modules proposed in DeepAAT, the following ablation experiments were conducted: 1⃝ The encoding order of GPS $L_{G}$ and the measurement matrix $L_{mes}$ is set to 2; 2⃝ The encoding order of GPS $L_{G}$ and the measurement matrix $L_{mes}$ is set to 4, which is the setting used in this article; 3⃝ The encoding order of GPS $L_{G}$ and the measurement matrix $L_{mes}$ is set to 6; 4⃝ Remove the spatial spectral feature aggregation module, similar to ESFM, using only the measurement matrix $\mathbf{W}_{mes}$ as the network input; 5⃝ Remove the global consistency-based outlier rejecting module. Here, the predicted pose of the network is directly used to triangulate all matching track points during global triangulation. These experimental data are the average of 10 test data results, which are shown in Tab.6. From Tab.6, it can be seen that:
Table 6: Results of ablation study
@@ 5.5 Ablation Study
(1) Encoding order $L$ of GPS information and the measurement matrix in the spatial-spectral feature aggregation module has little impact on the experimental results, indicating that as long as the encoding order is set within a reasonable range, good experimental results can be achieved.
@@
(3) Upon removal of the global consistency-based outlier rejecting module, there is an increase in the reprojection error. As shown in Fig.9, the noise feature points generated by triangulation increases significantly. The outliers included in the scene can have a negative impact on subsequent BA and easily lead to a local optimum. For example, for scene 8_127, the final result after BA is shown in Fig.10. It can be seen that there is a camera (marked with a green dashed box) whose pose has been optimized erroneously due to the influence of outliers.
@@
5.6 Generalization experiments on different sizes of input images
The proposed network does not require a fixed number of images as input. Therefore, to test the generalization of the network in scenarios with inconsistent camera numbers compared to the training sample, we apply the model trained on scene from 100-130 images to predict scenes with 30-50 cameras and 400-430 cameras. The predicted results are shown in Tab.7 and Tab.8.
Table 7: Results of outlier rejecting for 100-130 camera models in different numbers of image prediction tasks
@@ 5.6 Generalization experiments on different sizes of input images
Table 8: Pose prediction results of the 100-130 camera trained model in different numbers of camera prediction tasks
@@
The prediction results demonstrate that the proposed DeepAAT is versatile, excelling not only in scenes with a similar number of images but also in scenarios with significantly more or fewer images. As indicated in Tab.7, the network trained on scenes with 100-130 images shows a decline in accuracy, precision, recall, and F1 score when applied to 30-50 image scenes. Conversely, its performance metrics improve in 400-430 image scenes. This improvement could be attributed to the longer average track length of matching points in scenes with more cameras, increasing the likelihood of points being classified as inliers and reducing false negatives. This explains the high recall of 0.993 in scenes with 400-430 cameras. Regarding prediction time, the network demonstrates a notable efficiency: despite a rapid increase in the number of cameras, the time required for network prediction increases only marginally. This efficiency is a significant advantage in practical applications. In traditional AAT algorithms, both incremental and global, time consumption escalates quickly with an increasing number of scene images, a trend particularly pronounced in incremental SfM. Thus, our network’s time-saving benefits become more pronounced with larger sets of scene images.
From Tab.8, we observe that the network, trained on scenes with 100-130 images, exhibits a slight increase in average reprojection error, position error, and rotation error while predicting scenes with 30-50 images and those with 400-430 images. However, in scenarios such as 1_420 and 5_404, the network’s predicted position errors surpass the initial position errors. Nevertheless, following global BA, all scenes achieve accurate reconstruction results, as illustrated in Fig.11. This outcome not only underscores the network’s effective scene initialization capabilities but also highlights its robust reconstruction prowess.
@@ 6 Conclusion
7 Acknowledgment
This study was jointly supported by the National Natural Science Foundation Project (No. 42201477, No. 42130105).
[…]